# PDMS elastic properties: influence of fabrication protocol and test method

### v1

@@
Abstract
Silicone rubbers have been integral to the rapid development of microfluidic devices and flexible electronics in recent decades. Polydimethylsiloxane, commonly abbreviated as PDMS and produced using the brand name Dow SYLGARD 184, is a ubiquitous such rubber that is renowned for its ease of manufacturing and biocompatibility. SYLGARD 184 is a soft polymer frequently used in the field of mechanics, microfluidics, and electronics given the straightforward ability to manipulate its material properties via tuning base to curing agent mix ratios and curing schedules. However, this tuning process largely occurs on a case by case basis via trial and error, without a comprehensive understanding of how composition, temperature, or mechanical test method my alter the measured elasticity. Here we compare the reported elastic response of PDMS SYLGARD 184 obtained using a range of mechanical testing methods and compositions to investigate the variability. We also include previously unreported experimental results obtained in our lab using Volume Controlled Cavitation Expansion (VCCE). Both testing method and cure schedule (time and temperature of curing) are found to strongly influence the measured elastic response resulting in shear moduli that range over 3 orders of magnitude for chemical compositions that are reported to be the same. While tension testing is a common test method in mechanical analysis, the wide ranging moduli for stiff PDMS and complete inability to measure softer PDMS compositions reveal the limits of this standard method. As researchers in the field of mechanics strive to quantify the properties of increasingly complex biological and composite materials, converging on a standardized measurement of PDMS is necessary, and is a necessary first step for the community.
keywords:
polydimethylsiloxane (PDMS) , stiffness , SYLGARD 184
journal: Journal to be determined
@@ Abstract
[inst1]organization=Department of Mechanical Engineering, Massachusetts Institute of Technology, city=Cambridge, postcode=02139, state=MA, country=USA
[…]
[inst2]organization=Department of Civil and Environmental Engineering, Massachusetts Institute of Technology, city=Cambridge, postcode=02139, state=MA, country=USA
{highlights}
The stiffness of PDMS reported across literature is inconsistent
Tension testing, a mechanical “best practice”, produces the widest variability
False equivalencies arise when directly comparing different methods and cure schedules
Cure schedules and testing protocols should be standardized in the field
@@ 1 Introduction
Polydimethylsiloxane (PDMS) is widely used in microfluidics [1, 2, 3], medical devices, electronics [4, 5, 6], and engineering mechanics. As a two-part material system, the stiffness of the resulting silicone polymer can be easily tuned to suit the needs of an application. PDMS will cure at room temperature, or can have the curing accelerated with an oven cure. This flexibility of composition and cure is particularly useful in the context of biological research where PDMS has been proposed as a scaffold for muscle growth [7], and for studying cellular motility [8], as a biological membrane mimic in microfluidics [9], in biohybrid devices [10, 11, 12], and for cardiomyocyte growth [13, 14].
The tunability of bulk elastic and surface properties awarded by PDMS (or often the common branded product Dow SYLGARD 184 [15]) has driven its popularity in the fields of mechanics and advanced materials. PDMS is often used during method development for measurements of the elastic, adhesion, and fracture properties of soft materials [16, 17, 18, 19, 20]. Because of its relative chemical stability, many researchers also rely on PDMS during the development of tough materials that can undergo large deformations before failure[21, 22, 23, 24, 25, 26, 27].
The same fabrication flexibility that makes PDMS applicable across a range of fields leads to a critical problem of mechanical reproducibility: the elastic moduli of PDMS reported in literature vary widely. Authors rarely use consistent preparation and cure conditions unless they are from the same research team. Teixeira et al. [28] note this in their 2021 review, where they consolidate 25 articles on a range of preparation techniques that have been reported for PDMS used to perform a range of goals from stiffness modification, to foaming, to incorporation or inhomegeneities such as nanoparticles, metals, or carbon nanotubes. When looking at stiffness alone as a property of interest, the soft (elastic modulus $<$1 MPa) nature of PDMS further lends itself to a wider subset of mechanical characterization methods than are available for stiffer engineered materials. While standardized dogbone tension, compression, and impact tests are the standard for metal samples, as a soft material, PDMS has been tested with not only tension (both ASME standard and non standard) and compression but also indentation, nanoindentation, needle induced cavitation rheology, volume controlled cavity expansion (VCCE), and simple shear.
@@
“Round Robbin” studies have been used to compare test methodologies and increase the “reliability and repeatability” of measurements in ceramics [29, 30], composites [31], adhesives [32], concrete [33], steel [34], and many other materials and industries. In their Versailles Advanced Materials and Standards report, Kubler et al. [29] articulate their motivation for conducting the study to determine if the Single-edge-V-Notched beam method for measuring fracture toughness is “user-friendly, reliable and… comparable to other recognised methods”. A similar effort in the context of PDMS elasticity characterization is, in the view of the authors, overdue. To this end, we embark on this study in an effort to understand the magnitude of the problem at hand. We present the scale of variability observed across literature and hope to motivate the community to converge on more consistent preparation, and testing methodologies to improve repeatability and reprehensibility for PDMS and across the field of soft material mechanics.
@@
This letter is organized as follows: we begin by describing how studies were chosen for inclusion in this review (Sec. 2.1) and provide a brief description of the mechanical testing methods used across the included studies (Sec. 2.2). We then present the results of our analysis (Sec. 3) and include new data for moduli obtained via VCCE (detailed inB). Finally, we conclude with remarks and cautions on stiffnesses reported in literature for PDMS (Sec. 4).
@@ 2.1 Data gathering
Articles surveyed for this review were found through key word searches and references lists of included articles. Keywords included “elasticity” or “modulus” or “stiffness” and “PDMS” or “polydimethylsiloxane” or “SYLGARD 184”, as well as related derivatives and searches were performed across Google Scholar and Web of Science from January - August 2023. Studies included in the results presented here were those with defined a) mix ratios for curing agent and base, b) cure time, and c) cure temperature. We restrict this study to Dow’s SYLGARD 184. Hence, studies that used GE RTV PDMS silicones were excluded (such as [35] and some results reported in [36]). Note that one of the most commonly cited studies on PDMS moduli (See Fig. A1 in Section A), Lötters [4] has not been included as the authors used ABCR PS851 PDMS. All included studies are summarized in Table 3 of C and includes new data collected by the authors as detailed in B. Studies with parameters listed as “variable” intentionally survey a large range of values and are therefore not listed in the table (though included where appropriate in corresponding figures).
Dow recommends a mixing ratio of 10 parts base to 1 part curing agent (crosslinker) on a weight to weight basis; we will abbreviate this as 10:1 or more generally wB:wC. Where the mix ratio was instead reported as a percent, we have converted to the X parts base:1 part curing agent format. All stiffnesses are reported here as shear modulus, $\mu$. Where the authors initially report the stiffness as an elastic or Young’s Modulus, $E$, incompressibility is assumed with a Poisson’s ratio $\nu=0.5$ [37] and $\mu=E/(2(1+\nu))=3E$. For studies investigating viscoelastic material parameters or high strain rate response, the quasistatic modulus is reported in the summary charts and where only an instantaneous modulus is reported, we denote it as $\mu_{0}$. For studies reporting a cure at room temperature, 25 C was used in plotting and comparison tables.
Figure 1: Common types of testing used to determine shear or elastic modulus, $\mu$ and $E$ for PDMS. Variable definitions are included in the text.
@@ 2.2 Test methodologies
Tension: Tension testing was undertaken using a commercial tension test machine in all cases except [7] which used a fixed applied weight, [38] that uses a dynamic testing with a modified split Hopkinson bar, and [8] that uses a custom stretching tool and load cell. Tracking the force F, and sample extension $\Delta L$, from initial length $L$, the material stiffness is calculated as $\sigma=E\epsilon$ in the linear regime where $\sigma=F/A$ and $\epsilon=\Delta L/L$ are the stress and strain, respectively for a sample cross sectional area, $A$. Clamp displacement is used in [39, 40, 21, 36] for tracking average sample extension, while [38] use particle tracking and digital image correlation, [41, 42, 8] use optical image capture and post process with a known calibration, and [7] use direct measurement. The ASTM Standard D412 for a “Standard Test Methods for Vulcanized Rubber and Thermoplastic Elastomers—Tension” [43] includes a set of defined sample geometries, ramp rates, and test conditions and was used by [40, 44, 39, 45] who all focus on only a 10:1 base to curing agent ratio.
@@
Compression: Unconfined compression using $\sigma=E\epsilon$ in the linear regime as with tension testing was employed by [46] to compare against nanoindentation, as well as by [47] in their comparison of moduli across curing agent ratios.
@@
Nanoindentation: With an indenter tip possessing a projected contact area $A$, and geometry described by $\beta$, the modulus of a material near a free surface can be determined from the slope of the pressure-displacement curve $S=dP/dh$ of an indentation. Calculating the reduced modulus $E_{r}$ based on the indenter geometry, then combining with the indenter modulus $E_{i}$ and Poisson’s ratio $\nu_{i}$ returns the sample modulus, $E$. Commercially available hardware was used by most authors, however [45] were the only authors to use the standard analysis method included with their tool, while all other authors reported specifics of their analysis: [19] used the Oliver Pharr method (not accounting for adhesion) to determine $E_{r}^{OP}$. To avoid the overestimate of modulus that is typical in polymers when using Oliver Pharr, Cheng et al.[48] instead used a Hertzian contact model, whereby
$E_{H}=\sqrt{\frac{S^{3}(1-\nu^{2})^{2}}{6R_{i}P_{max}}}$ (1)
@@
for a spherical indentor with radius of curvature $R_{i}$ and $\nu$ is assumed to be 0.5 for PDMS. Multiple authors acknowledge the significance of adhesion in nanoindentation measurements and employ the Johnson, Kendall, Roberts (JKR) model instead of [46, 49, 20], or in addition to [50, 51], Hertzian contact. However authors disagreed whether the Hertzian contact model returned a higher moduli than the JKR model [46] or a lower moduli [20].
@@
Volume Controlled Cavity Expansion (VCCE): Controlled injection of an incompressible working fluid at the tip of an injection needle is used in this method to expand a spherical cavity in the sample [52, 17]. The resisting pressure in the cavity, $P$, is simultaneously measured to obtain a nonlinear pressure-volume relationship, or equivalently $P=P(a_{c})$, where $a_{c}$ is the effective cavity radius. Comparison with theoretical prediction of elastic, spherically symmetric, cavity expansion in incompressible hyperelastic materials, namely
@@
then allows to determine material parameters. Where $\lambda$ is the circumferential stretch which varies from $a_{c}/A_{c}$ at the cavity wall to $1$ in the remote field, $W(\lambda)$ is the corresponding elastic potential energy, and $A_{c}$ is the initial defect size, which needs to be determined. This method can also be applied to probe viscoelastic properties [53].
Li et al.[54] use similar assumptions in a cylindrical geometry, finding a very good match between analytical and experimental results.
Cavitation Rheology (CR): Alternately, a second type of cavity expansion, cavitation rheology, aims to identify the peak value of the resisting pressure, $P^{*}$, and assumes that it is well represented by the theoretical cavitation limit which is then used to determine the material parameters. In contrast to VCCE, this approach does not require control of the injected volume (it typically uses gas as an injection medium). Milner [55, 56] and Yang [57] assume a neo-Hookean material model [^1] This result can be obtained directly from (2) at the limit $\lambda\to\infty$, with a neo-Hooken form of $W(\lambda)$, as shown in [52]., which implies $P^{*}=(5/2)\mu_{nH}$.
Simple Shear: Both [58] and [59] assume linear (small strain) deformations in simple shear. An applied shear force, $F$, induces a displacement, $\Delta$, and corresponding shear strain $\gamma=\Delta/h$. The shear modulus is then estimated as $\mu=\tau/{\gamma}$, where the shear stress is $\tau=F/A$ and $h$ and $A$ are sample height and area shown in Fig. 1.
Beam Bending: This method relies on connected a displacement of a beam with know geometry to an applied load. Armani [37] measured tip deflection, $\delta$, of a length $L$ cantilever beam that was displaced under its own weight ($w$ per unit length) and calculated modulus using $\delta=\frac{wL^{4}}{8EI}$ with $I$ describing the second moment of area. By contrast, Du[60] considered a point load at the tip of a low aspect ratio beam and added a Timoshenko term $\alpha\frac{FL2(1+\nu)}{EA}$ with the shear coefficient $\alpha$.
Membrane Deflection: A unique measurement method [61], uses optical profilometry to measure the membrane deflection of a nanometer thickness layer of spin coated PDMS on a silicone wafer. By pressurizing a circular portion of the membrane with radius $r$ and thickness $t$, the authors were able to connect deflection, $\delta$, and material modulus as

### v2

# Explaining the spread in measurement of PDMS elastic properties: influence of test method and curing protocol Hannah Varner,a and Tal Cohena,b,∗

@@
Explaining the spread in measurement of PDMS elastic properties: influence of test method and curing protocol
Hannah Varner,a and Tal Cohena,b,∗
[Uncaptioned image] Accuracy in the measurement of mechanical properties is essential for precision engineering and for the interrogation of composition-property relationships. Conventional methods of mechanical testing, such as uniaxial tension, compression, and nanoindentation, provide highly repeatable and reliable results for stiff materials, for which they were originally developed. However, when applied to the characterization of soft and biological materials, the same cannot be said, and the spread of reported properties of similar materials is vast. Polydimethylsiloxane (PDMS), commonly obtained from Dow as SYLGARD 184, is a ubiquitous such material, which has been integral to the rapid development of biocompatible microfluidic devices and flexible electronics in recent decades. However, reported shear moduli of this material range over 2 orders of magnitude for similar chemical compositions. Taking advantage of the increased mechanical scrutiny afforded to SYLGARD 184 in recent years, we combine both published and new experimental data obtained using 9 mechanical test methods. A statistical analysis then elucidates the significant bias induced by the test method itself, and distinguishes this bias from the influence of curing protocols on the mechanical properties. The goal of this work is thus two-fold: (i) it provides a quantitative understanding of the different factors that influence reported properties of this particular material, and (ii) it serves as a cautionary tale. As researchers in the field of mechanics strive to quantify the properties of increasingly complex soft and biological materials, converging on a standardized measurement of PDMS is a necessary first step.
footnotetext: a Department of Mechanical Engineering. Massachusetts Institute of Technology. Cambridge, MA.footnotetext: bDepartment of Civil and Environmental Engineering. Massachusetts Institute of Technology. Cambridge, MA.footnotetext:  Corresponding author. E-mail: talco@mit.edu. 77 Massachusetts Avenue. Cambridge, MA 02139.footnotetext: † Electronic Supplementary Information (ESI) available: Details of original data published here for the first time, Summary of all articles included across the meta-analysis, including interconnection mapping of citations. Additional details of regression analysis. See DOI: XXXXXX/
@@ 1 Introduction
Modern engineering of material systems relies on the testing and quantification of a material’s mechanical response. The tension, compression, indentation and impact testing methods that most engineers first encounter are all standard for metal samples but may not be appropriate for soft materials, which are significantly more challenging to test. Nonetheless, the use of soft materials is becoming increasingly prevalent in engineering, as researchers try to better understand and interface with the human body, and develop smaller and more compliant material structures and devices. Comparing measurements from a range of techniques, several studies1, 2, 3 report wide discrepancies in measured stiffness of soft biological materials. It is notable that despite well documented variations in biological tissue, these discrepancies are often attributed primarily to the testing method and not to material variations.
Even commercially available synthetic materials are hard to test. One example is Polydimethylsiloxane (PDMS) where reported elastic moduli range by 2 orders of magnitude for similar compositions, as shown herein. PDMS is a two-part material system and the stiffness of the resulting silicone polymer can be easily tuned to suit the needs of an application. PDMS will cure at room temperature, or can have the curing accelerated with an oven cure. It is widely used in microfluidics 4, 5, 6, medical devices, and electronics 7, 8, 9. In engineering mechanics, it is a common surrogate material for method development in the measurement of elastic, adhesion, and fracture response 10, 11, 12, 13, 14, and as a component in the development of tough materials that can undergo large deformations before failure15, 16, 17, 18, 19, 20, 21. The flexibility of composition and cure is particularly useful in the context of biological research where PDMS has been proposed as a scaffold for cell growth 22, 23, 24, 25, as a biological membrane mimic in microfluidics 26, and in biohybrid devices 27, 28, 29. Given its widespread use in biology, numerous authors have worked to modify the adhesion, wetablity, absorptivity, dielectric properties, and mechanical characteristics of PDMS, with comprehensive reviews provided by Wolf et al. 30, Zhou et al.31, Abbasi et al. 32, Teixeira et al. 33 and Murphy et al. 34. However, far less attention has been paid to systematically understanding and measuring the bulk properties of PDMS (or, often, the commercially branded product Dow SYLGARD 184 35). The same fabrication flexibility and wide availability that makes commercial SYLGARD 184 applicable across a range of fields also highlights a critical problem of mechanical reproducibility.
The primary goal of this work is to elucidate the bias induced by the testing method in determining the elastic response of soft materials. To this end, we choose SYLGARD 184 as a test case, taking advantage of the breadth of available data in the literature spanning a range of test methods, and combined with new Volume Controlled Cavity Expansion (VCCE) data, reported here. The secondary goal of this work is, thus, the interrogation of SYLGARD 184 mechanical properties. While frequent users of SYLGARD 184 often expect minor inconsistencies in the mechanical properties due to changes from the manufacturer***The exact composition of SYLGARD 184 is proprietary. Reading the product literature and safety data sheets provides the following insights 30, 36, 37, 38: The base contains a majority dimethylvinyl-terminated dimethyl siloxane ($>$60%), with a significant portion of silica fillers (30-60% dimethylvinylated and trimethylated silica), as well as 1-5% tetra(trimethylsiloxy) silane and small amounts of ethylbenzene, xylene and a platinum catalyst (¡1% each). The curing agent is a majority the cross linking agent dimethyl methylhydrogen siloxane (40-70%), additional methylated silica and siloxanes, and 1-5% the inhibitor tetramethyl tetravinyl cyclotetrasiloxane, we hypothesize instate that measurement discrepancies can be mainly attributed to bias imparted by: (1) different mechanical testing methods, and (2) wide variation in the cure conditions used for similar mixing ratios.
@@
“Round Robin” studies have been used to compare test methodologies and increase the “reliability and repeatability” of measurements in ceramics 39, 40, composites 41, adhesives 42, concrete 43, steel 44, and many other materials and industries. In their Versailles Advanced Materials and Standards report, Kubler et al. 39 articulate their motivation for conducting the study to determine if the Single-edge-V-Notched beam method for measuring fracture toughness is “user-friendly, reliable and … comparable to other recognised methods”. A similar effort in the context of PDMS elasticity characterization is, in the view of the authors, overdue. To this end, we embark on this study in an effort to understand the magnitude of the problem at hand. We present the scale of variability observed across literature and in our own data, and hope to motivate the community to converge on more consistent preparation, and testing methodologies to improve repeatability for PDMS and across the field of soft material mechanics.
@@
This paper is organized as follows: we begin by describing how studies were chosen for inclusion in this meta-analysis (Sec. 2.1) and provide a brief description of the mechanical testing methods used across the included studies (Sec. 2.2). We then present the results of our analysis (Sec. 3) and include new data for moduli obtained via VCCE (detailed in Sec. 2 of the ESI, Table ESI.T1). Finally, we conclude with remarks and cautions on stiffnesses reported in literature for PDMS (Sec. 4).
@@ 2.1 Data gathering
Articles surveyed for this review were found through key word searches and references lists of included articles. Keywords included “elasticity” or “modulus” or “stiffness” and “PDMS” or “polydimethylsiloxane” or “SYLGARD 184”, as well as related derivatives and searches were performed across Google Scholar and Web of Science from January - August 2023. Studies included in the results presented here were those with defined a) mix ratios for curing agent and base, b) cure time, and c) cure temperature. We restrict this study to Dow’s SYLGARD 184. Hence, studies that used other commercial formulations such as GE RTV PDMS silicones were excluded (such as by Liu et al.45 and some results reported by Schneider et al. 46). Note that one of the most commonly cited studies on PDMS moduli (see Fig. ESI.F1 in ESI.1), Lötters 7 has not been included as the authors used ABCR PS851 PDMS.
All included studies are summarized in Table ESI.T2 of Sec. ESI 3. Studies with parameters listed as “variable” intentionally survey a large range of values and are therefore not listed in the table (though included where appropriate in corresponding figures). New data collected by the authors and included in Table ESI.T2, is described in Sec. ESI.2 with fabrication and measurement detailed in ESI.T1.
Dow recommends a mixing ratio of 10 parts base to 1 part curing agent (crosslinker) on a weight to weight basis; we will abbreviate this as 10:1 or more generally wB:wC. Where the mix ratio was instead reported as a percent, we have converted to the X parts base:1 part curing agent format. All stiffnesses are reported here as shear modulus, $\mu$. Where authors initially reported the stiffness as an elastic or Young’s Modulus, $E$, incompressibility is assumed with a Poisson’s ratio $\nu=0.5$ 47 and $\mu=E/(2(1+\nu))=3E$. For studies investigating viscoelastic material parameters or high strain rate responses, the quasistatic modulus is reported in the summary charts and where only an instantaneous modulus is reported, we denote it as $\mu_{0}$. For studies reporting a cure at room temperature, 25 C was used in plotting and comparison tables.
Fig.  1: Common types of testing used to determine shear or elastic modulus, $\mu$ and $E$ for PDMS. Variable definitions are included in the text.
@@ 2.2 Test methodologies
Tension: Tension testing was undertaken using a commercial tension test machine in all cases except Brown et al. 22 who use a fixed applied weight, Upadhyay et al.48 who perform dynamic testing with a modified split Hopkinson bar, and Fuard et al. 25 who use a custom stretching tool and load cell. Tracking the force $F$, and sample extension $\Delta L$, from initial length $L$, the material stiffness is calculated as $\sigma=E\epsilon$ in the linear regime where $\sigma=F/A$ and $\epsilon=\Delta L/L$ are the stress and strain, respectively, for a sample cross sectional area, $A$. Clamp displacement is used for tracking average sample extension in 49, 50, 51, 52, 15, 46, while Upadhyay et al. 48 use particle tracking and digital image correlation, Mills et al. 53, Seghir and Arscott 54 and Fuard et al. 25 use optical image capture and post process with a known calibration, and Brown et al. 22 use direct measurement. The ASTM Standard D412 for a “Standard Test Methods for Vulcanized Rubber and Thermoplastic Elastomers—Tension” 55 includes a set of defined sample geometries, ramp rates, and test conditions and was used by 52, 49, 56, 51, 57 who all focus on only a 10:1 base to curing agent ratio.
@@
Compression: Unconfined compression (using $\sigma=E\epsilon$ to determine the elastic stiffness in the linear regime as with tension testing) was employed by Carrillo et al. 58 to compare against nanoindentation, as well as by Wang et al. 59 in their comparison of moduli across curing agent ratios.
@@
Nanoindentation: With an indenter tip possessing a projected contact area $A$, and geometry described by $\beta$, the modulus of a material near a free surface can be determined from the slope of the pressure-displacement curve $S=dP/dh$ of an indentation. Calculating the reduced modulus $E_{r}$ based on the indenter geometry, then combining with the indenter modulus $E_{i}$ and Poisson’s ratio $\nu_{i}$ returns the sample modulus, $E$. Commercially available hardware was used by most authors, however, Mata et al.57 were the only authors to use the standard analysis method included with their tool, while all other authors reported specifics of their analysis: Patel et al.13 used the Oliver Pharr method (not accounting for adhesion) to determine $E_{r}^{OP}$. To avoid the overestimate of modulus that is typical in polymers when using Oliver Pharr, Cheng et al.60 instead used a Hertzian contact model, whereby
$E_{H}=\sqrt{\frac{S^{3}(1-\nu^{2})^{2}}{6R_{i}P_{max}}}$
@@
for a spherical indenter with radius of curvature $R_{i}$ and $\nu$ is assumed to be 0.5 for PDMS. Multiple authors acknowledge the significance of adhesion in nanoindentation measurements and employ the Johnson, Kendall, Roberts (JKR) model instead of Hertzian contact 58, 61, 14, or in addition to it62, 63. However, authors disagreed whether the Hertzian contact model returned a higher moduli than the JKR model 58 or a lower moduli 14.
@@
Volume Controlled Cavity Expansion (VCCE): Controlled injection of an incompressible working fluid at the tip of an injection needle is used to expand a spherical cavity in a sample 64, 11. The resisting pressure in the cavity, $P$, is simultaneously measured to obtain a nonlinear pressure-volume relationship, or equivalently $P=P(a_{c})$, where $a_{c}$ is the effective cavity radius. Comparison with theoretical prediction of elastic, spherically symmetric, cavity expansion in incompressible hyperelastic materials, namely
@@
then allows to determine material parameters. Where $\lambda$ is the circumferential stretch which varies from $a_{c}/A_{c}$ at the cavity wall to $1$ in the remote field, $W(\lambda)$ is the corresponding elastic potential energy, and $A_{c}$ is the initial defect size, which needs to be determined. This method can also be applied to probe viscoelastic properties 65.
Li et al.66 use similar assumptions in a cylindrical geometry, finding a very good match between analytical and experimental results.
Cavitation Rheology (CR): Alternately, a second type of cavity expansion, cavitation rheology, aims to identify the peak value of the resisting pressure, $P^{*}$, and assumes that it is well represented by the theoretical cavitation limit which is then used to determine the material parameters. In contrast to VCCE, this approach does not require control of the injected volume (it typically uses gas as an injection medium). Milner 67, 68 and Yang et al. 69 assume a neo-Hookean material modelThis result can be obtained directly from Eq. (1) at the limit $\lambda\to\infty$, with a neo-Hookean form of $W(\lambda)$, as shown in 64., which implies $P^{*}=\frac{5}{2}\mu_{nH}$.
Simple Shear: Both Upadhyay et al.70 and Nunes 71 assume linear (small strain) deformations in simple shear. An applied shear force, $F$, induces a displacement, $\Delta$, and corresponding shear strain $\gamma=\Delta/h$. The shear modulus is then estimated as $\mu=\tau/{\gamma}$, where the shear stress is $\tau=F/A$ and $h$ and $A$ are sample height and area shown in Fig. 1.
Beam Bending: This method relies on connecting the displacement of a beam with know geometry to an applied load. Armani et al. 47 measured tip deflection, $\delta$, of a length $L$ cantilever beam that was displaced under its own weight ($w$ per unit length) and calculated modulus using $\delta=\frac{wL^{4}}{8EI}$ with $I$ describing the second moment of area. By contrast, Du et al.72 considered a point load at the tip of a low aspect ratio beam and added a Timoshenko term $\frac{\alpha FL}{EA}$ with the shear coefficient $\alpha$.
Membrane Deflection: A unique measurement method by Thangawng et al. 73, uses optical profilometry to measure the membrane deflection of a nanometer thickness layer of spin coated PDMS on a silicone wafer. By pressurizing a circular portion of the membrane with radius $r$ and thickness $t$, the authors were able to connect deflection, $\delta$, and material modulus as